# Feasible Space Monitoring for Multiple Control Barrier Functions with application to Large Scale Indoor Navigation

### v1

@@ Abstract
Quadratic programs (QP) subject to multiple time-dependent control barrier function (CBF) based constraints have been used to design safety-critical controllers. However, ensuring the existence of a solution at all times to the QP subject to multiple CBF constraints is non-trivial. We quantify the feasible solution space of the QP in terms of its volume. We introduce a novel feasible space volume monitoring control barrier function that promotes compatibility of barrier functions and, hence, existence of a solution at all times. We show empirically that our approach not only enhances feasibility but also exhibits reduced sensitivity to changes in the hyperparameters such as gains of nominal controller. Finally, paired with a global planner, we evaluate our controller for navigation among humans in the AWS Hospital gazebo environment. The proposed controller is demonstrated to outperform the standard CBF-QP controller in maintaining feasibility.
@@ I INTRODUCTION
Videos accompanying examples and simulation results as well as a short description of our work can be seen at https://www.youtube.com/watch?v=N94NqPKjRSM.
@@
In this work, we analyze the conditions under which a controller with multiple CBF constraints admit a solution for all time, a property we call persistent feasibility. At a given state and time, we quantify the feasible control input space defined by CBF constraints and input bounds in terms of its volume, and introduce a novel feasible-space control barrier function (FS-CBF) that restricts the rate-of-change of the volume of the feasible space. For control-affine dynamics and polytopic input constraints, this feasible space takes the form of a convex polytope. Existence of an FS-CBF is shown to be a necessary and sufficient condition for enforcing compatibility of multiple CBFs.
@@
Two commonly used approaches are employed in the literature for handling multiple CBF constraints to tackle the feasibility problem: combining all barrier functions into a single one [1, 2, 3] or imposing the barriers simultaneously [4, 5, 6]. In [1, 3], barrier functions are combined into a single barrier constraint through a smoothed minimum (or maximum) operator. Employing non-smooth analysis, [2] constructs a barrier function with the non-smooth minimum (or maximum) operation. When multiple CBF constraints are imposed in a CBF-QP formulation, [4, 5] assumes compatibility and hence existence of a suitable control input. A constraint relaxation scheme is proposed in [6] whenever CBF-QP becomes infeasible, albeit without guarantees of safety.
@@
A few studies explicitly try to circumvent the persistent feasibility issue. A sampling-and-grid refinement method is proposed in [7] to search the domain of state space over which multiple control barrier functions are compatible. However, it does not propose a controller to confine the robot to this domain.
@@
Figure 1: Indoor navigation scenario. (a) Robot (green circle) navigating in AWS Hospital gazebo environment using proposed CBF controller. (b) Rviz visualization of the global map, global planner’s path (green), humans (blue) and the nearest static obstacles (red, found in the way described in Section VI-B) used by the robot for collision avoidance.
@@
A sum-of-squares (SoS)-based CBF synthesis method when constraint sets are described by polynomial functions is given in [8]. In [9], sufficient conditions are provided for designing multiple CBFs concurrently so that a QP-based controller using all the CBFs will always be feasible. The work by [10] consolidates multiple CBFs into one CBF and proposes a predictor-corrector adaptation law for weights of constituent constraint functions. Additionally, a CBF learning framework is proposed in [11] for decentralized but cooperative multi-agent systems. Some works allow class-$\mathcal{K}$ functions to adapt by having time-varying parameters. The work by [12] creates a dataset of persistently feasible and eventually infeasible parameters by randomly sampling parameters and simulating the system forward. Adaptation of class-$\mathcal{K}$ parameters is also performed by posing CBF-QP as a differentiable layer in a predictor-corrector framework [13] and in learning paradigm [14] for learning of safe policies. Our motivation lies in ensuring safety in highly constrained indoor environments, such as in Fig. 1 with static as well as dynamically moving agents like humans. The decentralized nature of this problem and resulting large state space leads to prohibitive amount of possible scenarios thereby precluding synthesis of a valid CBFs akin to most of existing works.
With our FS-CBF, we provide a method to enforce compatibility between different CBF conditions. In cases where synthesis of our CBF is computationally feasible using any available algorithm, the existence of FS-CBF may be seen as complementing the existing literature. In other cases, we show through simulations that imposing feasible space volume as a candidate FS-CBF enhances persistent feasibility of standard CBF formulations, and additionally reduces sensitivity of CBF-QP controllers to hyperparameters such as gains of the nominal controller. This work, to the best of our knowledge, is the first to consider an optimization problem’s feasible solution space volume and its gradient in control design. Previous approaches simply check if the CBF-QP is feasible, which amounts to checking whether the volume is non-zero or not. Finally, paired with ROS2 navigation stack planners, we also provide the first application of CBF-QP controller on a large scale indoor navigation scenario by implementing our controller on AWS Hospital environment in Gazebo [15].
@@
  • Introduced the notion of persistent feasibility and formulated Feasible-Space Control Barrier Function (FS-CBF) to quantify and restrict the rate-of-change of the volume of the feasible control input space.
  • Demonstrated through simulations that FS-CBF enhances persistent feasibility and reduces hyperparameter sensitivity in CBF-QP controllers in a large scale indoor navigation scenario, integrated with ROS2 navigation stack planners.
@@ III System Description
(Control Barrier Function) [16, 17] For the dynamical system (1), $h_{i}:\mathbb{R}^{+}\times\mathcal{X}\rightarrow\mathbb{R}$ is a control barrier function (CBF) on the set $\mathcal{S}_{i}(t)$ defined by (2a)-(2c) if there exists an extended class-$\mathcal{K}$ function $\alpha_{i}:\mathbb{R}\rightarrow\mathbb{R}^{+}$ and a set $\mathcal{K}_{i}$ with $\mathcal{S}_{i}\subseteq\mathcal{K}_{i}\subset\mathbb{R}^{n}$ such that $\forall x\in\mathcal{K}_{i},\forall t>0$
$\sup_{u\in\mathcal{U}}\left[\frac{\partial h_{i}(t,x)}{\partial t}+L_{f}h_{i}(t,x)+L_{g}h_{i}(t,x)u\right]\geq-\alpha_{i}(h_{i}(t,x))$ (3)
It is assumed in this work that each $h_{i}$ is a CBF for the set $\mathcal{S}_{i}$ and that there exists a Lipschitz continuous controller $u:\mathbb{R}^{+}\times\mathcal{S}_{i}\rightarrow\mathcal{U}$ that satisfies (3), guaranteeing thus the forward invariance of the set $\mathcal{S}_{i}$ under the closed-loop dynamics (1) [16, 17]. Finally, we review the following QP controller that is often used to enforce multiple CBF constraints
$u(t,x)=\arg\min\quad$ $||u-u_{ref}(t,x)||^{2}$ (4)
s.t. $\dot{h}_{i}(t,x)\geq\alpha_{i}(h_{i}(t,x)),i\in\{1,..,N\},$
@@ IV Problem Formulation
$\mathcal{U}_{i}(t,x)=\{u\in\mathbb{R}^{m}~{}|~{}\dot{h}_{i}(t,x,u)+\alpha_{i}(h_{i}(t,x))\geq 0\}.$ (5)
@@
Feasible CBF Space: At a given time $t$ and state $x\in\mathcal{X}$, the feasible CBF space $\mathcal{U}_{c}\subset\mathbb{R}^{m}$ is the intersection of aforementioned sets with control input domain
@@
For control-affine dynamics, the constraints (5) are affine in $u$ and represent halfspaces. If the input domain is polytopic, that is, $\mathcal{U}=\{u\in\mathbb{R}^{m}~{}|~{}A_{u}u\leq b_{u}\},A_{u}\in\mathbb{R}^{2m\times m},b\in\mathbb{R}^{2m}$, then the feasible space $\mathcal{U}_{c}$ is a convex polytope. Figure 2 visualizes the polytope for a single integrator-modeled robot as it navigates around obstacles. For brevity, we will be denoting the feasible space in following form:
@@
where $A:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{(N+2m)\times m},b:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{(N+2m)\times 1}$. We also denote each row of $A$ by $a_{i}$ and each element of $b$ by $b_{i},i\in\{1,2,..,N\}$.
Figure 2: Feasible space (green) of a single integrator agent with X, Y velocity inputs as it navigates to its goal location with N=3 in (4) while avoiding obstacles (black). (a) The path in configuration space, (b), (c), (d) feasible control space visualization at different times. The dotted lines represent hyperplane equations that constrain the control input. The $2\times 2$ square is formed by control input bound. Other hyperplanes are formed as a result of CBF equations.
@@
(Persistently Feasible CBFs) Given a Lipschitz continuous controller $\pi:\mathcal{X}\rightarrow\mathcal{U}$, the functions $h_{i}$ are persistently feasible for dynamics (1) with $u=\pi(x)$ in (1) if $\exists E\subset\mathcal{X}$ such that $\mathcal{U}_{c}(t,x)\neq\emptyset$ $\forall t\geq 0,x_{0}\in E$.
We now describe our objective for persistent feasibility.
@@
Consider the dynamical system (1) subject to $N$ barrier function constraints $h_{i}(t,x)\geq 0$ where each $h_{i}$, defined in (III), is a CBF with class-$\mathcal{K}$ function $\alpha_{i}$ for the set $\mathcal{S}_{i}$ respectively. If for the initial state $x_{0}\in\mathbb{R}^{n}$, $\mathcal{U}_{c}(0,x_{0})\neq\emptyset$, design a controller $\pi$ that renders the CBFs $h_{i},i\in\{1,..,N\}$ persistently feasible, that is, $\mathcal{U}_{c}(t,x)\neq 0,\quad\forall t>0$.
[Uncaptioned image] Figure 3: (a) Configuration space of green robot with illustrative orange obstacles, (b) The feasible control space $\mathcal{U}_{c}$ defined by multiple CBFs at the current state and time. It’s volume is given by the area of the green region. (c) The state space domain (example) over which CBFs are compatible, i.e., volume of the polytope $\mathcal{U}_{c}$ is $>0$. The controlled invariant set is a subset of domain of compatibility that can actually be maintained forward invariant.
[Uncaptioned image] Figure 4: Feasible space (green) of H1-H4 and its inscribing circle. H5 (H6) intersects the feasible space (circle) at a single point.
@@ V Feasible Volume Control Barrier Function
The infeasibility of CBF-QP results when $\mathcal{U}(t,x)$ becomes empty. In this section, we propose a novel control barrier function built upon the feasible space as a means to solving Problem 1. First, we provide some preliminaries on computing exact and approximate volume of convex polytopes and then present our formal results on FS-CBF.
V-A Polytope Volume Estimation
@@
V-A1 Monte Carlo
A common approach to estimating polytope volume is to evaluate the integral with Monte Carlo (MC) sampling[18, 19]. This method is known to work well for low-dimensional spaces, but not for high-dimensional spaces[20]. Moreover, since Monte Carlo estimates are prone to be noisy, this method is also not suitable for evaluating gradients, which are needed to develop CBF conditions.
Below we describe two other approximate ways to compute the feasible space volume. It is worth mentioning that 2,3 are under-approximations, while MC can over- or under-approximate the feasible space volume depending on the samples.
V-A2 Chebyshev Ball of a Polytope

### v2

@@ Abstract
Quadratic programs (QP) subject to multiple time-dependent control barrier function (CBF) based constraints have been used to design safety-critical controllers. However, ensuring the existence of a solution at all times to the QP subject to multiple CBF constraints (hereby called compatibility) is non-trivial. We quantify the feasible control input space defined by multiple CBFs at a state in terms of its volume. We then introduce a novel feasible space (FS) CBF that prevents this volume from going to zero. FS-CBF is shown to be a sufficient condition for the compatibility of multiple CBFs. For high-dimensional systems though, finding a valid FS-CBF may be difficult due to the limitations of existing computational hardware or theoretical approaches. In such cases, we show empirically that imposing the feasible space volume as a candidate FS-CBF not only enhances feasibility but also exhibits reduced sensitivity to changes in the user-chosen parameters such as gains of the nominal controller. Finally, paired with a global planner, we evaluate our controller for navigation among other dynamically moving agents in the AWS Hospital gazebo environment. The proposed controller is demonstrated to outperform the standard CBF-QP controller in maintaining feasibility.
@@ I INTRODUCTION
In this work, we introduce a CBF that enforces the existence of a common solution to multiple CBF constraints at all times, a property we call persistent compatibility. Towards this, we introduce the following new metric. At a given state and time, we quantify the feasible control input space defined by multiple CBF constraints and input bounds in terms of its volume. For example, for control affine dynamics and polytopic input constraints, the feasible control space is also given by a convex polytope. Next, we introduce a novel feasible-space control barrier function (FS-CBF) that prevents this volume from going to zero. Our controller is thus a QP that augments the user-provided CBF constraints with an additional FS-CBF constraint. The existence of FS-CBF is shown to be a sufficient condition for enforcing the persistent compatibility of multiple CBFs.
@@
Two commonly used approaches employed in the literature for handling multiple CBF constraints are: combining all barrier functions into a single one [1, 2, 3] or imposing the barriers simultaneously [4, 5, 6]. In [1, 3], barrier functions are combined into a single barrier constraint through a smoothed minimum (or maximum) operator. Employing non-smooth analysis, [2] constructs a barrier function with the non-smooth minimum (or maximum) operation. All of the aforementioned approaches assume compatibility of CBFs when imposed together [4, 5] or assume that the newly constructed CBF is a valid CBF for dynamics of the system and input bounds.
@@
A few studies explicitly try to circumvent the persistent compatibility issue. A sampling-and-grid refinement method is proposed in [7] to search the domain of state space over which multiple control barrier functions are compatible. However, it does not propose a controller to confine the robot to this domain.
@@
Figure 1: Indoor navigation scenario (a) Robot (green circle) navigating in AWS Hospital Gazebo environment using proposed CBF controller. (b) Rviz visualization of the global map, global planner’s path (green), humans (blue) and the nearest static obstacles (red, found in the way described in Section VI-C) used by the robot for collision avoidance.
@@
A sum-of-squares (SoS)-based CBF synthesis method when constraint sets are described by polynomial functions is given in [8]. In [9], sufficient conditions are provided for designing multiple CBFs concurrently so that a QP-based controller using all the CBFs will always be feasible. The work by [10] consolidates multiple CBFs into one CBF and proposes a predictor-corrector adaptation law for weights of constituent constraint functions. An offline-sampling procedure is proposed in [11] to construct a CBF from constraint functions when the former is related to the latter through a constant scaling and offset of state vector. Additionally, a CBF learning framework is proposed in [12] for decentralized but cooperative multi-agent systems. Some works allow class-$\mathcal{K}$ functions to adapt by having time-varying parameters. The work by [13] creates a dataset of persistently compatible and eventually incompatible parameters by randomly sampling parameters and simulating the system forward. Adaptation of class-$\mathcal{K}$ parameters is also performed by posing CBF-QP as a differentiable layer in a predictor-corrector framework [14] and in learning paradigm [15] for learning of safe policies.
We introduce a procedure that exploits existing approaches such as reachability analysis [16] to synthesize a valid FS-CBF. Owing to the limitations of these approaches though, it might not always be computationally feasible to find FS-CBF. In such cases, we show empirically that that imposing feasible space volume as a candidate FS-CBF enhances the persistent compatibility of standard CBF formulations. It is also observed to reduce the sensitivity of CBF-QP controllers to user-chosen parameters such as gains of the nominal controller. Furthermore, finding the volume of feasible space and its gradient with respect to constituent hyperplanes are non-trivial problems as they do not accept analytical expressions in general. As a contribution, we also introduce some of the existing approximations suitable for control-oriented applications. To the best of our knowledge, we are the first to present this paradigm of quantifying the feasible control input space of CBFs as well as consider an optimization problem’s feasible solution space volume and its gradient in control design. Previous approaches simply check if the CBF-QP is feasible, which amounts to checking whether the volume is non-zero. Note that instead of synthesizing FS-CBF, it may be possible to consolidate multiple CBFs into a single valid CBF offline. However, such an ideal scenario is difficult, if not impossible, to achieve due to prohibitive computational requirements of existing approaches. It may also be difficult to account for all possible possibilities in the offline phase. For example, such scenarios may have varying number of humans and obstacles. Our motivation lies in ensuring safety in highly constrained indoor environments, such as in Fig. 1 with static as well as dynamically moving agents like humans. Our idea to monitor feasible space volume thus has a high practical appeal for robotic applications.
Finally, as a case study, paired with ROS2 navigation stack planners, we also provide the application of CBF-QP controller on a large scale indoor navigation scenario by implementing our controller on AWS Hospital environment in Gazebo [17]. Several existing approaches such as [18, 19, 20] for indoor navigation with humans rely only on fast replanning through a low-level planner based on simplified dynamics for collision avoidance, and do not employ a safety-critical controller. Our controller may thus complement existing planning frameworks. Among CBF-based approaches, [21] uses CBF-MPC, an MPC controller with constraints defined by CBF derivative conditions, for navigation in crowd, and [22] designs a novel CBF for polytopic obstacles and uses it in the CBF-MPC framework for navigation in corridor environments. From an application perspective, [21] does not consider indoor environments and [22] does not consider dynamic moving agents. While MPC does help with maintaining feasibility of CBF constraints, it is unlike our method of monitoring the volume of feasible space.
@@
  • Introduced the notion of persistent compatibility and formulated Feasible-Space Control Barrier Function (FS-CBF) to quantify and restrict the rate-of-change of the volume of the feasible control input space.
  • Demonstrated through simulations that FS-CBF enhances persistent feasibility and reduces hyperparameter sensitivity in CBF-QP controllers. We also demonstrate our controller in a large-scale indoor navigation scenario, integrated with ROS2 navigation stack planners.
@@ III System Description
When the relative degree $r_{i}$ of $h_{i}$ is greater than one, we can derive the following new barrier functions
$\psi^{k}_{i}(t,x)=\dot{\psi}^{k-1}_{i}(t,x)+\alpha_{i}^{k}\psi_{i}^{k}(t,x)$ (3)
for $k=\{1,..,r_{i}-1\}$, $\psi_{i}^{0}=h_{i}$ and where $\alpha_{i}^{k}$ are extended class-$\mathcal{K}$ functions.
@@
(Higher Order CBF)[23, 24] The function $h_{i}(t,x):\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a Higher-Order CBF (HOCBF) of $r_{i}$-th order on the set $\mathcal{S}_{i}(t)$ if there exist $r_{i}$ extended $\mathcal{K}$ functions $\alpha_{i}^{k}:\mathbb{R}\rightarrow\mathbb{R},\;k\in\{1,2,..,r_{i}\}$, and an open set $D_{i}(t)\subset\mathbb{R}^{+}\times\mathbb{R}^{n}$ with $\mathcal{S}_{i}(t)\subset D_{i}(t)\subset\mathcal{X}$ such that
$\dot{\psi}_{i}^{r_{i}-1}(t,x,u)\geq-\alpha_{i}^{r_{i}}(\psi_{i}^{r_{i}-1}(t,x)),~{}\forall x\in D_{i}(t),\forall t\geq 0.$ (4)
Note that for $r_{i}=1$, (4) reduces to $\dot{h}_{i}(x,u)\geq-\alpha_{i}h_{i}(x)$. And for $r_{i}>1$, $\dot{\psi}_{i}^{k}$ does not depend on $u$ for $k<r_{i}-1$. It is assumed in this work that each $h_{i}$ is a HOCBF for the set $\mathcal{S}_{i}$ and that there exists a Lipschitz continuous controller $u:\mathbb{R}^{+}\times\mathcal{S}_{i}\rightarrow\mathcal{U}$ that satisfies (4), guaranteeing thus the forward invariance of the set $\mathcal{S}_{i}$ under the closed-loop dynamics (1) [25, 26]. Finally, we review the following controller that is often used to enforce multiple CBF constraints
$u(t,x)=\arg\min_{{\normalsize{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}u\in\mathcal{U}}}}}\quad$ $\!||u-u_{ref}(t,x)||^{2}$ (5)
s.t. $\dot{\psi}_{i}^{r_{i}-1}(t,x,u)\geq\alpha_{i}^{r_{i}}(\psi_{i}^{r_{i}-1}(t,x))$
$\quad\quad\quad\quad\quad\quad\quad\quad i\in\{1,..,N\}$
@@ IV Problem Formulation
$\mathcal{U}_{i}(t,x)=\{u\in\mathbb{R}^{m}~{}|~{}\dot{\psi}_{i}^{r_{i}-1}(t,x,u)+\alpha_{i}^{r_{i}}(\psi_{i}^{r_{i}-1}(t,x))\geq 0\}.$ (6)
@@
Feasible CBF Space: At a given time $t$ and state $x\in\mathcal{X}$, the feasible CBF space $\mathcal{U}_{c}\subset\mathbb{R}^{m}$ is the intersection of sets $\mathcal{U}_{i},i\in\{1,..,N\}$ with control input domain
@@
For control-affine dynamics, the constraints (6) are affine in $u$ and represent halfspaces. Note that for control-affine dynamics, (4) and (6) are affine in the control input $u$, and therefore (5) is a QP controller when $\mathcal{U}=\mathbb{R}^{m}$. If the input domain is also polytopic, that is, $\mathcal{U}=\{u\in\mathbb{R}^{m}~{}|~{}A_{u}u\leq b_{u}\},A_{u}\in\mathbb{R}^{2m\times m},b\in\mathbb{R}^{2m}$, then the feasible space $\mathcal{U}_{c}$ is a convex polytope, and the controller (5) is still a QP, since only linear constraints have been added. Figure 2 visualizes the polytope for a single integrator-modeled robot as it navigates around obstacles. For brevity, we will be denoting the feasible space in the following form:
@@
where $A:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{(N+2m)\times m},b:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{(N+2m)\times 1}$. We also denote each row of $A$ by $a_{i}$ and each element of $b$ by $b_{i},i\in\{1,2,..,N\}$. Next, we formally introduce the notion of compatibility and state the objective of this paper.
Figure 2: Feasible space (green) of a single integrator agent with X, Y velocity inputs (in m/s) as it navigates to its goal location with N=3 in (5) while avoiding obstacles (black). (a) The path in configuration space, (b), (c), (d) feasible control space visualization at different times. The dotted lines represent hyperplane equations that constrain the control input. The $2\times 2$ square is formed by control input bound. Other hyperplanes are formed as a result of CBF equations.
@@
(Compatible CBFs) At time $t$ and state $x$, the barriers $h_{i},i\in\{1,..,N\}$ are called compatible if $\mathcal{U}_{c}(t,x)\neq\emptyset$. Furthermore, given a Lipschitz continuous controller $\pi:\mathcal{X}\rightarrow\mathcal{U}$, the barriers $h_{i}$ are called persistently compatible for dynamics (1) with $u=\pi(x)$ in (1) if $\exists E\subset\mathcal{X}$ such that $x_{0}\in E\implies\mathcal{U}_{c}(t,x)\neq\emptyset$ for all $t\geq 0$.
@@
Consider the dynamical system (1) subject to $N$ barrier function constraints $h_{i}(t,x)\geq 0$ where each $h_{i}$, defined in (2), is a CBF with class-$\mathcal{K}$ function $\alpha_{i}$ for the set $\mathcal{S}_{i}$ respectively. If for the initial state $x_{0}\in\mathbb{R}^{n}$, $h_{i}(t,x_{0})>0,i\in\{1,..,N\},~{}\mathcal{U}_{c}(0,x_{0})\neq\emptyset$, design a controller such that $\forall t>0,~{}h_{i}(t,x(t))>0$ and $\mathcal{U}_{c}(t,x)>0$.
We propose a solution to Problem 1 in Section V-A. However, owing to limitations of related literature, our solution may be difficult to realize in practice. We therefore propose several approximations in Sections V-B and V-C and provide empirical validation of their efficacy in Sections VI-B and VI-C.
@@ V Feasible Volume Control Barrier Function
The infeasibility of CBF-QP results when $\mathcal{U}(t,x)$ becomes empty. In this section, we propose a novel control barrier function built upon the feasible space as a means to solving Problem 1.
V-A Polytope Volume Barrier Function
A measure of the feasible space is its volume. Let $\mathcal{V}(t,x)=\textrm{vol}(\mathcal{U}_{c})(t,x)$ where vol is the volume function that, for a polytope $P=\{u|Au\leq b\}$, is defined as follows
@@ V-A Polytope Volume Barrier Function
where $du_{j}$ is an infinitesimal change in $u_{j}$. Note that $\mathcal{V}(t,x)=0$ when $\mathcal{U}_{c}(t,x)=\emptyset$. The compatible state space is defined as the set of states $D(t)=\{x\in\mathcal{X}~{}|~{}\mathcal{V}(t,x)\geq 0\}$ and is visualized in Fig. 3. We then aim to find a CBF $\mathcal{V}_{c}:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$, with class-$\mathcal{K}$ function $\alpha_{\mathcal{V}_{c}}$, that renders $D(t)$ forward invariant by imposing the following condition
$\dot{\mathcal{V}_{c}}(t,x,u)\geq-\alpha_{\mathcal{V}_{c}}(\mathcal{V}_{c}(t,x))$ (10)
We now present the following theorem that states the sufficiency of FS-CBF condition (10) for compatibility of $h_{i}$.
Theorem 1.
Let $h_{i},i\in\{1,..,N\}$ be $N$ barrier functions corresponding to $N$ safe sets defined as in (2). Let $\mathcal{V}(t,x)$ be the volume of the feasible CBF space $\mathcal{U}_{c}(t,x)$ in the control domain defined in (7). For $10<\epsilon<<$ 1, let $D(t)=\{x\in\mathcal{X}~{}|~{}\mathcal{V}(t,x)\geq\epsilon\}$ be the compatible state space of $h_{i},i\in\{1,..,N\}$. Then $h_{i}$ are persistently compatible if there exists $\mathcal{C}\subset\mathcal{D}$ and a function $\mathcal{V}_{c}:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that $\mathcal{V}_{c}$ is a control barrier function on +×$\mathbb{R}^{+}\times\mathcal{C}$ and $x(0)\in\mathcal{C}$.
Proof.
Suppose that $\mathcal{V}_{c}$ is a CBF with class-$\mathcal{K}$ function $\alpha_{\mathcal{V}}$ on +×$\mathbb{R}^{+}\times\mathcal{C}$, with $\mathcal{C}\subset\mathcal{D}$. If $x(0)\in\mathcal{C}$, then $x(t)\in\mathcal{C},\forall t>0$ under the action of the controller that satisfies (10). Therefore, $\mathcal{V}_{c}(t,x(t))\geq 0,\forall t>0$ and hence $\mathcal{V}(t,x(t))\geq\epsilon$ for all $t>0$, implying persistent compatibility. ∎
Note that sets of measure 0 have zero volume. Therefore, in Theorem V, we enforce $\mathcal{V}\geq\epsilon$ to distinguish between the case when $\mathcal{U}_{c}=\emptyset$ from the case when $\mathcal{U}_{c}$ is a set of measure 0 (for example, a line in 3-dimensional space) as the latter does not exclude compatibility. Also, note that (9) implies that when $\mathcal{U}_{c}=\emptyset$, $\mathcal{U}{(t,x)}=0$ and therefore $\mathcal{V}$ is well-defined in the entire state space.
Computing $\mathcal{V}(t,x)$ requires computing the volume of the feasible space $\mathcal{U}_{c}$. This is not a trivial problem and even when feasible space is shaped as a convex polytope as in (8), it does not admit analytical expressions for $\mathcal{V}$. In the next section, we introduce some existing approximate approaches that are suitable for our application.
Figure 3: (a) Workspace of black robot with illustrative orange obstacles, (b) The state space. In the violet region, CBFs are compatible, i.e., the volume of the polytope $\mathcal{U}_{c}$ is $>0$. The yellow region is the largest control invariant set contained in the violet region. (c) The feasible control space $\mathcal{U}_{c}$ defined by multiple CBFs at the current state and time. It’s volume is given by the area of the green region in this 2D example.
Figure 4: Feasible space (green) defined by hyperplanes H1-H5. H3, H4 are overlapping. Suppose H4, H5 are free to move horizontally. It can be shown that the gradient of the area of the green region exists w.r.t position of H5 but not w.r.t H4.
Contrary to intuition, only imposing (10) in a controller does not ensure that the chosen control input lies inside the feasible space. Since instantaneously violating a CBF constraint does not lead to an immediate violation of safety, choosing a control input outside $\mathcal{U}_{c}(t,x)$ may help in increasing $\mathcal{V}_{c}$ and therefore satisfies the condition (10). $\mathcal{V}_{c}$ is strictly used to enforce compatibility of $h_{i}$. To ensure that chosen control input maintains compatibility of $h_{i}$ as well as lies inside the feasible space $\mathcal{U}_{c}$, we design the following controller
$\pi(t,x)\!=\!\arg\min_{u\in\mathcal{U}}\quad$ $||u-u_{ref}(t,x)||^{2}+M\delta^{2}$ (11a)
s.t. $\dot{\mathcal{V}_{c}}(t,x,u)\!\geq-\alpha_{\mathcal{V}_{c}}(\mathcal{V}_{c}(t,x))+\delta$ (11b)
$\psi_{i}^{r_{i}-1}(t,x,u)\geq\alpha_{i}^{r_{i}}(\psi_{i}^{r_{i}-1}(t,x))$
$\quad\quad\quad\quad\quad\quad\quad\quad i\in\{1,..,N\}$ (11c)
where $\delta$ is a slack variable used to relax (11b) and $M>>1$. The relaxation is necessary as maintaining compatibility of CBF constraints is not equivalent to ensuring safety. To see this, note that CBF constraints are mathematically well-defined even when the state is unsafe, that is, $x\notin S_{i}~{}\forall i$, and act like a Lyapunov function to attract the state $x$ to the safe set. Therefore, since choosing a control input that satisfies (11c) would, loosely speaking, ensure safety (see Remark 1), and (11b) does not have a one-to-one relationship with safety, (11b) and (11c) need not have a common solution. To prioritize safety, we enforce (10) along with (4) in (11) but with a slack variable. In theory, therefore, because of $\delta$ (11) does not possess guarantees of compatibility and could lead to infeasibility. Finding conditions under which compatibility and safety can be maintained simultaneously is subject to future work. This work, nonetheless provides a novel perspective on CBF-QP controllers and in Section VI, we focus on empirical showing improvements over standard CBF-QP (5).
Remark 1.
Note that (5) and (11) may not result in Lipschitz-continuous control inputs. To the best of our knowledge, obtaining a holistic solution that guarantees smoothness and uniqueness of solutions and ensures safety along with compatibility of constraints for (1) is an open problem. Further investigation of safety in presence of discontinuous control input, similar to [4], is subject to future work.
Remark 2.
We demonstrate in Section VI-B and Fig. 6 that imposing constraint (11) results in $u$ being chosen in the interior of $\mathcal{U}_{c}$, even when the nominal control input is outside the feasible space. This is done to prevent the volume from shrinking too fast. The controller (5) on the hand always chooses a point on the boundary of feasible space. A mapping from $u_{ref}$ to the interior of feasible space was also designed in [27] through a gauge map. The gauge map was designed in a different way than ours and focused on obtaining a differentiable control design with single barrier constraint only.
V-B Polytope Volume as Candidate CBF
In Section VI-A, we show an application of an existing CBF synthesis method to construct $\mathcal{V}_{c}$ for a simple system. For highly dynamic and constrained environments though, as previously mentioned, it is usually computationally prohibitive to consider all possible scenarios in the design of CBFs. Therefore, as is typically done in practice, we assume the constraint function is the CBF, that is, $\mathcal{V}_{c}=\mathcal{V}$ and use the condition (10) to promote persistent compatibility. When imposing $\mathcal{V}_{c}=\mathcal{V}$, we impose linear class-$\mathcal{K}$ functions in our simulation and experiment results. We also relax the constraint (11b) with a slack variable
V-C Polytope Volume Estimation
In this section, we describe several ways to approximately compute the volume of feasible space.
V-C1 Monte Carlo
A common approach to estimating polytope volume is to evaluate the integral with Monte Carlo (MC) sampling[28, 29]. We randomly sample $K$ particles $p_{i}\in\mathcal{U}$ and evaluate the volume as follows
$\mathcal{V}(t,x)=\frac{\sum_{j=1}^{K}\mathds{1}(p_{i}\in\mathcal{U}_{c})}{\textrm{vol}(\mathcal{U})}$ (12)
where $\textrm{vol}(\mathcal{U})$ is assumed to be given and easy to compute practically as it is often defined as a symmetric polygon.
V-C2 Chebyshev Ball of a Polytope